# Mutual Information Surprise: Rethinking Unexpectedness in Autonomous Systems

### v1

@@ 3.1 What Do We Expect from a Surprise?
where $\hat{I}_{n}$ is the estimation of mutual information $I_{n}$ at the time of the first $n$ observations, and $\hat{I}_{n+m}$ for $I_{m+n}$ after observing $m$ additional points. Starting from here, we omit the variable terms, $\mathbf{x}$ and $\mathbf{y}$, in the notations of mutual information and its estimation for the sake of simplicity. A large (relative to sample size $m$ and $n$) positive MIS signals enlightenment, indicating significant learning, whereas a near-zero or negative MIS indicates frustration, suggesting stalled progress. Hence, MIS provides operational insight into whether a system evolves as expected, transforming it into a practical autonomy test. Significant deviations from the expected MIS trajectory indicate meaningful changes or system stagnation.
@@ 3.2 Bounding MIS
Mutual information estimation is inherently challenging: it is high-dimensional, nonlinear, and exhibits complex variance. The standard method, though principled, is a computationally expensive permutation test  (?, ?), involving repeatedly shuffling $m+n$ observations into two groups, calculating MI differences, and evaluating rejection probabilities:
@@
    The number of existing observations $n$ is much smaller than cardinality of space $\mathcal{X},\mathcal{Y}$. $n\ll|\mathcal{X}|,|\mathcal{Y}|$
@@
Consider a well-regulated autonomous system defined in Section 3.1, which satisfies the conditions in Assumption 1. With probability at least $1-\rho$, the change in MLE-based mutual information estimates satisfies:
@@
The proof of Theorem 1 is shown in the Appendix. These bounds are both tighter ($\mathcal{O}(\frac{\log n}{n})$ instead of $\mathcal{O}(\frac{\log n}{\sqrt{n}})$) and more efficient (analytical test sequence) than previous methods. The bounds offer theoretically grounded thresholds within which we expect MI to evolve. When these bounds $MIS_{\pm}$ are breached—either from below or from above—we know the system has encountered something.
Some may argue that for an oversampled system, Assumption 2 does not hold. That is true, and as a result, the expectation term in Theorem 1, $\log(m+n)-\log n$, needs to be adjusted. For a noise-free system with limited outcome and large number of existing observations, one needs to replace the expectation term with $(|\mathcal{Y}|-1)(\frac{1}{n}-\frac{1}{m+n})$ and the bounds in Theorem 1 still works.
@@ 3.3 What Does MIS Actually Tell Us?
When the quantity $\text{MIS}=\hat{I}_{n+m}-\hat{I}_{n}$ falls outside the established bounds $MIS_{\pm}$—either exceeding the upper bound or falling below the lower bound—the system is considered to be surprised, thereby triggering a Mutual Information Surprise (MIS). Essentially, Theorem 1 functions as a statistical hypothesis test: the null hypothesis posits that the underlying system remains well-regulated, implying $\Delta I=I_{n+m}-I_{n}=0$, where $I_{n}$ denotes the true mutual information at the time of $n$ observations. Any violation indicates a significant shift, with negative deviations ($\Delta I<0$) and positive deviations ($\Delta I>0$) each carrying distinct implications.
@@
this implies $\Delta I(\mathbf{x},\mathbf{y})<0$, signifying a downward shift in mutual information. A negative surprise indicates diminished or stalled learning, potentially due to:
@@
this implies $\Delta I(\mathbf{x},\mathbf{y})>0$, indicating an upward shift in mutual information. This positive surprise can result from:
@@
Summary Table
@@
Violation from Below Stagnation in exploration $\downarrow H(\mathbf{x})\Rightarrow\downarrow I(\mathbf{x},\mathbf{y})$
Increased noise / process drift $\uparrow H(\mathbf{y}\mid\mathbf{x})\Rightarrow\downarrow I(\mathbf{x},\mathbf{y})$
Violation from Above Aggressive exploration $\uparrow H(\mathbf{x})\Rightarrow\uparrow I(\mathbf{x},\mathbf{y})$
Noise reduction $\downarrow H(\mathbf{y}\mid\mathbf{x})\Rightarrow\uparrow I(\mathbf{x},\mathbf{y})$
Novel discovery $\uparrow H(\mathbf{y})\Rightarrow\uparrow I(\mathbf{x},\mathbf{y})$
The table above summarizes potential causes for MIS violations and their implications. These patterns help the system differentiate between meaningful learning and misleading deviations, expanding beyond the capacity of classical surprise measures and providing a road map to corrective or adaptive responses for higher level autonomy. We purposely omit the case where a decrease in $H(\mathbf{y})$ causes violation from below, as this scenario typically lacks independent significance. Instead, its happening is generally caused by changes in sampling strategy or underlying processes, which we have already discussed.
@@ 3.4 Reaction Policy: A Three-Pronged Approach
2. Process Forking. The second policy responds to variations in conditional entropy $H(\mathbf{y}\mid\mathbf{x})$, i.e., changes in function mapping. Upon surprise triggered by $\Delta\hat{H}(\mathbf{y}\mid\mathbf{x})$, the system forks into two subprocesses, each consisting of $n$ existing observations and $m$ new observations divided at the surprise moment (Theorem 1). The two subprocesses represent the prior process (existing observations) and the likely altered process (new observations), and will continue their sampling separately. The subprocess first encountering a $\Delta\hat{H}(\mathbf{y}\mid\mathbf{x})$-triggered surprise is discarded, and the remaining subprocess continues as the main process. In the extremely rare case when both subprocesses trigger a $\Delta\hat{H}(\mathbf{y}\mid\mathbf{x})$ dominated MIS surprise at the same time, we discard the process with fewer observations, and continues with the subprocess with more observations.
@@ 4.1 Putting Surprise to the Test
The surprise measures are computed as follows. Shannon Surprise is calculated using its classical definition in Eq. (1), as the negative log-likelihood of the true label under a Gaussian Process predictive model. Bayesian Surprise is computed as Postdictive Surprise, defined in Eq. (2), using the KL divergence between the prior and posterior predictive distributions of $y$ at each input $x$. The same Gaussian Process predictive model is used for both, using a Matérn $\nu=2.5$ kernel with a constant noise level set to $0.1$. After each surprise computation, the model is re-trained with all currently available data.
@@
Scenario 1: Standard Exploration.
New data is randomly sampled from $x\in[30,100]$, expanding the domain without altering the underlying function or aggressively exploring unfamiliar regions. This represents a system exploring new yet consistent areas of its environment.
Expected behavior: A well-calibrated surprise measure should indicate ongoing learning without abrupt fluctuations. We do not expect MIS to be violated.
As shown in Figure 1, MIS progresses steadily within its expected bounds, reflecting a stable and well-regulated learning process. In contrast, Shannon and Bayesian Surprises fluctuate erratically, often spiking without clear justification.
@@
Scenario 2: Over-Exploitation.
In this scenario, the system repeatedly samples a previously seen point from $x\in[0,30]$, specifically observing the pair $(x,y)=(7,7)$ one hundred times. This simulates stagnation.
@@
Figure 2 shows that MIS falls below its lower bound, signaling a lack of knowledge gain. While Shannon and Bayesian Surprises also trend downward, they lack a defined lower threshold, limiting their reliability for flagging such behavior. Recall that both Shannon and Bayesian Surprises are inherently one-sided, as noted in  (?) and  (?).
Figure 2: Surprise measures under over-exploitation.
Scenario 3: Noisy Exploration.
We perform standard exploration over $x\in[30,100]$ but apply random corruption to the outputs $\mathbf{y}$, replacing each with a uniformly random digit between 0 and $9$. This simulates exploration without informative feedback.
@@
Figure 3 confirms the following: MIS drops below its expected range, accurately signaling lack of learning. In contrast, Shannon and Bayesian Surprises again display erratic behavior without consistent trends.
Figure 3: Surprise measures under noisy exploration.
Scenario 4: Aggressive Exploration.
This scenario enforces strict exploration over $x\in[30,500]$, where each new sample is far from all observed points (i.e., outside the $\pm 1$ neighborhood range).
@@
Figure 4 shows MIS surpassing its upper bound, consistent with this expectation. Shannon and Bayesian Surprises again fluctuate unpredictably.
Figure 4: Surprise measures during aggressive exploration.
Scenario 5: Noise Decrease.
To simulate noise reduction, we begin with $100$ initial observations from $x\in[0,30]$, paired with a randomly assigned output $y\in[0,9]$. New samples are drawn from the same $x$ range but the new $y$ is produced using the deterministic modulus function in Eq. (7).
@@
Figure 5 confirms this: MIS grows beyond its bound. Shannon and Bayesian Surprises continue to spike erratically.
Figure 5: Surprise measures during noise decrease.
Scenario 6: Discovery of New Output Values.
We modify the function in the unexplored region ($x>30$) to $y=x\mod 10-10$, introducing a different behavior while keeping the original function unchanged in $[0,30]$.
@@
Across all scenarios, MIS reliably indicates whether the system is genuinely learning, stagnating, or encountering degradation. It responds to the structure and value of observations, more than just novelty. In contrast, Shannon and Bayesian Surprises often react to superficial fluctuations and display numerical instability. Furthermore, the MIS progression bound remains consistent and interpretable across all scenarios, while Shannon and Bayesian Surprises lack a universal scale or threshold, as reflected by their inconsistent magnitudes across Figures 1 through 6. This inconsistency limits their effectiveness as a reliable trigger. Overall, this simulation study demonstrates MIS not only as a novel metric for quantifying surprise, but also as a more trustworthy indicator of learning dynamics—making it a promising tool for autonomous system monitoring.
@@ 4.2 Pollution Estimation: A Case Study
As discussed in Section 3.4, the MIS reaction policy is designed to complement existing exploration–exploitation strategies. To demonstrate the effectiveness of the Mutual Information Surprise Reaction Policy (MISRP), we integrate it with three well-established sampling strategies. These are: the surprise-reactive (SR) sampling method proposed by  (?) using either Shannon or Bayesian surprises, the subtractive clustering/entropy (SC/E) active learning strategy proposed by  (?), and the greedy search/query by committee (GS/QBC) active learning strategy used in  (?).
@@
Shannon and Bayesian Surprise are computed following the procedure described in Section 4.1. For MIS calculations, we discretize the range of pollution values observed in the data into $100$ bins to estimate entropy.
@@
Across all comparisons, the baseline strategies display considerable volatility. In contrast, MISRP-governed counterparts produce smoother and consistently lower error curves, highlighting the stabilizing effect of MIS through its ability to facilitate adaptive responses in dynamic environments.
Table 2 presents the average estimation errors and their corresponding standard errors. The standard error is measured across $10$ Monte Carlo simulations and $450$ frames. Across all sampling strategies, incorporating the MIS reaction policy yields a substantial reduction in both mean estimation error and variability. Improvements in estimation error range from $24\%$ to $76\%$, while reductions in standard error range from $36\%$ to $90\%$.
@@
Sampling Strategy Estimation Error (Vanilla) Estimation Error (MISRP-Governed) Mean Improvement Std. Error Improvement
SR with Shannon $6.64\pm 0.436$ $\mathbf{1.60\pm 0.043}$ $76\%$ $90\%$
SR with Bayesian $2.79\pm 0.096$ $\mathbf{0.87\pm 0.016}$ $69\%$ $83\%$
SC/E $2.02\pm 0.071$ $\mathbf{1.53\pm 0.045}$ $24\%$ $36\%$
GS/QBC $2.07\pm 0.071$ $\mathbf{1.49\pm 0.039}$ $28\%$ $45\%$
Table 2: Comparison of pollution map estimation errors: baseline sampling strategies versus MISRP-governed strategies.

### v3

@@ 3.1 An Alternative Surprise Definition
where $\hat{I}_{n}$ denotes the estimate of mutual information based on the first $n$ observations, and $\hat{I}_{n+m}$ denotes the estimate after observing an additional $m$ samples, with $\mathbf{x}$ and $\mathbf{y}$ omitted for simplicity. In this work, we adopt the standard estimation approach of histogram-based distribution representation (?) and a default of $10$ bins unless otherwise specified.
A large positive MIS indicates enlightenment (increase in $I(\mathbf{x};\mathbf{y})$), meaning that the new observations significantly improve the system understanding of the input-output relationship. In contrast, a near-zero or negative MIS indicates frustration (decrease in $I(\mathbf{x};\mathbf{y})$), suggesting that new data contributes little to learning or even disrupts previously acquired structure. In this way, MIS provides a practical signal for assessing whether an autonomous system continues to acquire knowledge from its environment. We argue that MIS is a better metric for guiding decision making of autonomous systems and we will provide further support to sustain our claim in the subsequent sections.
@@ 3.2 Bounding MIS
Testing the change in $I(\mathbf{x};\mathbf{y})$ via MIS is challenging: mutual information estimation is nonlinear and exhibits complex variance. The standard method, though principled, is a computationally expensive permutation test (?, ?), involving repeatedly shuffling $m+n$ observations into two groups, calculating MI differences, and evaluating rejection probabilities:
@@
    The number of existing observations $n$ is much smaller than the cardinality of space $\mathcal{X},\mathcal{Y}$. $n\ll|\mathcal{X}|,|\mathcal{Y}|$
@@
Consider a well-regulated autonomous system defined in Definition 1, which satisfies the conditions in Assumption 1. With probability at least $1-\rho$, the change in MLE-based mutual information estimates satisfies:
@@
The proof of Theorem 1 is shown in the Appendix. These bounds are both tighter ($\mathcal{O}(\frac{\log n}{n})$ instead of $\mathcal{O}(\frac{\log n}{\sqrt{n}})$) and more efficient (analytical test sequence) than previous methods. The bounds offer theoretically grounded thresholds within which we expect MI to evolve. When these bounds $MIS_{\pm}$ are breached—either from below or from above—we then deem that the true mutual information of the system has changed.
@@ 3.3 What Does MIS Actually Tell Us?
When the quantity $\text{MIS}=\hat{I}_{n+m}-\hat{I}_{n}$ falls outside the established bounds $MIS_{\pm}$—either exceeding the upper bound or falling below the lower bound—the system is considered to be surprised, thereby triggering a Mutual Information Surprise (MIS). Essentially, Theorem 1 functions as a statistical hypothesis test: the null hypothesis posits that the underlying system remains well-regulated, implying $\Delta I=I_{n+m}-I_{n}=0$, where $I_{n}$ denotes the true mutual information at the time of $n$ observations. Any violation indicates a significant shift, with negative deviations ($\Delta I<0$) and positive deviations ($\Delta I>0$) each carrying distinct implications on learning performance via the Fano Lemma (?).
@@
this implies $\Delta I(\mathbf{x};\mathbf{y})<0$, signifying a downward shift in mutual information. A negative surprise indicates diminished or stalled learning, potentially due to:
@@
this implies $\Delta I(\mathbf{x};\mathbf{y})>0$, indicating an upward shift in mutual information. This positive surprise can result from:
@@
Table 1: Summary Table: MIS Violations and Their Potential Causes
@@
Violation from Below Stagnation in exploration $\downarrow H(\mathbf{x})\Rightarrow\downarrow I(\mathbf{x};\mathbf{y})$
Increased noise / process drift $\uparrow H(\mathbf{y}\mid\mathbf{x})\Rightarrow\downarrow I(\mathbf{x};\mathbf{y})$
Violation from Above Aggressive exploration $\uparrow H(\mathbf{x})\Rightarrow\uparrow I(\mathbf{x};\mathbf{y})$
Noise reduction $\downarrow H(\mathbf{y}\mid\mathbf{x})\Rightarrow\uparrow I(\mathbf{x};\mathbf{y})$
Novel discovery $\uparrow H(\mathbf{y})\Rightarrow\uparrow I(\mathbf{x};\mathbf{y})$
Table 1 summarizes potential causes for MIS violations and their implications. These patterns help the system differentiate between meaningful learning and misleading deviations, expanding beyond the capacity of classical surprise measures and providing a road map to corrective or adaptive responses for higher level autonomy. We purposely omit the case where a decrease in $H(\mathbf{y})$ causes violation from below, as this scenario typically lacks independent significance. Instead, its happening is generally caused by changes in sampling strategy or underlying processes, which we have already discussed.
@@ 3.4 Reaction Policy: A Three-Pronged Approach
2. Process Forking. The second policy responds to variations in conditional entropy $H(\mathbf{y}\mid\mathbf{x})$, i.e., changes in conditional distribution. Upon surprise triggered by $\Delta\hat{H}(\mathbf{y}\mid\mathbf{x})$, the system forks into two subprocesses, each consisting of $n$ existing observations and $m$ new observations divided at the surprise moment (Theorem 1). The two subprocesses represent the prior process (existing observations) and the likely altered process (new observations), and will continue their sampling separately. The subprocess first encountering a $\Delta\hat{H}(\mathbf{y}\mid\mathbf{x})$-triggered surprise is discarded, and the remaining subprocess continues as the main process. In the extremely rare case when both subprocesses trigger a $\Delta\hat{H}(\mathbf{y}\mid\mathbf{x})$ dominated MIS surprise at the same time, we discard the process with fewer observations, and continues with the subprocess with more observations.
@@ 4.1 Putting Surprise to the Test
The surprise measures are computed as follows. Shannon Surprise is calculated using its classical definition in Eq. (3), namely as the negative log-likelihood of the true label under a Gaussian Process (GP) predictive model. Bayesian Surprise is computed using Postdictive Surprise, defined in Eq. (6), by evaluating the KL divergence between the prior and posterior predictive distributions of $\mathbf{y}$ at each input $\mathbf{x}$.
While the current definition of Shannon and Bayesian Surprises are single instance, they can be extended to multi-instance settings. Specifically, for a sequence of $m$ independently sampled observation pairs $(\mathbf{X},\mathbf{Y})=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{m}$, the cumulative Shannon Surprise (CSS) is defined as
$S_{\text{Shannon}}(\mathbf{X},\mathbf{Y})=\sum_{i=1}^{m}S_{\text{Shannon}}(\mathbf{x}_{i},\mathbf{y}_{i}\mid p_{i-1}),$ (12)
where $p_{i-1}$ denotes the predictive distribution $p(\mathbf{y}\mid\mathbf{x})$ prior to observing $(\mathbf{x}_{i},\mathbf{y}_{i})$. This formulation is also known as the cumulative prequential log score (?), which has been widely used in sequential concept drift and anomaly detection literature (?, ?, ?).
For Bayesian (Postdictive) Surprise, we define the multi-instance version as the divergence between predictive distributions evaluated on the most recent observation $(\mathbf{x}_{m},\mathbf{y}_{m})$, after updating parameters from $\boldsymbol{\theta}$ to $\boldsymbol{\theta}^{\prime}$ using the dataset $(\mathbf{X},\mathbf{Y})$. Formally,
$S_{\text{Postdictive}}(\mathbf{X},\mathbf{Y})=D_{\text{KL}}\left(P(\mathbf{y}_{m}\mid\boldsymbol{\theta}^{\prime},\mathbf{x}_{m})\,\|\,P(\mathbf{y}_{m}\mid\boldsymbol{\theta},\mathbf{x}_{m})\right).$
We consider two practical multi-instance variants: a cumulative version that aggregates surprise values from the beginning of the sequence, and a rolling-window version that measures cumulative surprise within the most recent $W=20$ sampling steps.
All versions of Shannon and Bayesian Surprises use the same Gaussian Process predictive model, with a Matérn kernel ($\nu=2.5$) and a fixed noise level of $0.1$. Except for the multi-instance Bayesian Surprise, the model is retrained using all available observations after each surprise computation.
@@
Scenario 1: Standard Exploration. New data is randomly sampled from $x\in[30,100]$, expanding the domain while adhering to the same underlying response function in Eq. (11). As the sampling region broadens, the input distribution $p(\mathbf{x})$ evolves, which in turn alters the joint distribution $p(\mathbf{x},\mathbf{y})$ and the observed functional relationship $p(\mathbf{y}\mid\mathbf{x})$. Notably, this observed mapping may differ from the true system mapping in sparsely explored regions, where it is effectively inferred through extrapolation rather than direct observation.
This setting exemplifies a well-regulated autonomous system as defined in Definition 1: although the joint distribution and the observed mapping evolve over time, the underlying input-output relationship, and thus the mutual information $I(\mathbf{x};\mathbf{y})$, remains stable.
Expected behavior: Since the system is well-regulated, a meaningful surprise measure should not react to exploration-driven distributional shifts. In particular, variations in $p(\mathbf{x})$ and the induced changes in $p(\mathbf{y}\mid\mathbf{x})$ should not be interpreted as unexpected events. Accordingly, we do not expect MIS to be violated.
As shown in Figure 1, MIS evolves steadily within its expected bounds, reflecting stable learning despite continuous changes in the observed data distribution. In contrast, the single-instance Shannon and Bayesian Surprises fluctuate erratically, frequently producing spikes that lack clear interpretability. Their cumulative and rolling variants, while smooth, tend to exhibit largely monotonic trends that offer little actionable insight for decision-making. Notably, the cumulative Shannon Surprise shows a pivot around $25$ sampling steps, which can be attributed to increased coverage of the sampling space: as the model becomes more familiar with the environment, familiar observations begin to dominate and outweigh unseen observations in new samples.
Figure 2: Surprise measures under over-exploitation.
Scenario 2: Over-Exploitation. In this scenario, the system repeatedly samples a previously seen point from $x\in[0,30]$, specifically observing the pair $(x,y)=(7,7)$ one hundred times. This simulates stagnation.
@@
Figure 2 shows that MIS falls below its lower bound, clearly signaling a lack of knowledge gain. Although single-instance Shannon and Bayesian Surprises also exhibit a downward trend, they lack a principled lower threshold, limiting their ability to reliably detect such behavior. As noted in (?) and (?), both Shannon and Bayesian Surprises are inherently one-sided measures, which further constrains their interpretability. Similar patterns are observed for their multi-instance variants, with one notable exception: cumulative Bayesian Surprise increasingly labels the exploitation behavior as more surprising over time, whereas rolling Bayesian measures deem it progressively less surprising. Nevertheless, due to its KL-divergence formulation, Bayesian Surprise lacks a clear probabilistic threshold, rendering this trend difficult to translate into actionable signals.
Figure 3: Surprise measures under noisy exploration.
Scenario 3: Noisy Exploration. We perform standard exploration over $x\in[30,100]$ but apply random corruption to the outputs $\mathbf{y}$, replacing each with a uniformly random digit between 0 and $9$. This simulates exploration without informative feedback.
@@
Figure 3 confirms this pattern: MIS drops below its expected range, accurately signaling knowledge loss. In contrast, Shannon and Bayesian Surprises again exhibit erratic behavior without consistent trends. Meanwhile, the multi-instance Shannon Surprise diverges to infinity, and the cumulative and rolling variants of Bayesian Surprise display contradictory monotonic patterns, further limiting their interpretability and practical usefulness.
Figure 4: Surprise measures during aggressive exploration.
Scenario 4: Aggressive Exploration. This scenario enforces strict exploration over $x\in[30,500]$, where each new sample is far from all observed points (i.e., outside the $\pm 1$ neighborhood range).
@@
Figure 4 shows MIS exceeding its upper bound, consistent with the expected behavior under pure exploration. In contrast, the single-instance Shannon and Bayesian Surprises again fluctuate unpredictably. The multi-instance Shannon Surprise diverges to infinite values, while the cumulative and rolling versions of Bayesian Surprise once more exhibit contradictory monotonic trends, limiting their interpretability.
Figure 5: Surprise measures during noise decrease.
Scenario 5: Noise Decrease. To simulate noise reduction, we begin with $100$ initial observations from $x\in[0,30]$, paired with a randomly assigned output $y\in[0,9]$. New samples are drawn from the same $x$ range but the new $y$ is produced using the deterministic modulus function in Eq. (11).
@@
Figure 5 confirms this: MIS grows beyond its bound. Shannon and Bayesian Surprises shows similar behaviors to the prior scenarios.
Scenario 6: Discovery of New Output Values. We modify the function in the unexplored region ($x>30$) to $y=x\mod 10-10$, introducing a different behavior while keeping the original function unchanged in $[0,30]$.
@@
Across all scenarios, MIS reliably indicates whether the system is genuinely learning, stagnating, or encountering degradation. It responds to the structure and value of observations, more than just novelty. In contrast, single-instance Shannon and Bayesian Surprises often react to superficial fluctuations and display numerical instability, and multi-instance Shannon and Bayesian Surprises often exhibits simple and even contradictory (cumulative versus rolling) monotone behaviors. Furthermore, the MIS progression bound remains consistent and interpretable across all scenarios, while Shannon and Bayesian Surprises lack a universal scale or threshold, as reflected by their inconsistent magnitudes across Figures 1 through 6. This inconsistency limits their effectiveness as a reliable trigger. Overall, this simulation study demonstrates MIS not only as a novel metric for quantifying surprise, but also as a more trustworthy indicator of learning dynamics—making it a promising tool for autonomous system monitoring.
Table 2 summarizes the differences between Mutual Information Surprise vs Shannon and Bayesian family of surprises.
@@ 4.2 Pollution Estimation: A Case Study
As discussed in Section 3.4, the MIS reaction policy (MISRP) is designed as a data-stream management mechanism that complements existing exploration-exploitation strategies. To evaluate its effectiveness, we conduct simulations using three well-established sampling strategies: the surprise-reactive (SR) sampling method proposed by (?), implemented with either Shannon or Bayesian Surprises; the subtractive clustering/entropy (SC/E) active learning strategy introduced by (?); and the greedy search/query-by-committee (GS/QBC) active learning strategy used in (?). In addition, input-output data stream management has long been studied in the concept drift and anomaly detection literature. Accordingly, we compare MISRP against one of the most widely used cumulative drift detection metrics, the cumulative Shannon Surprise (CSS) defined in Eq. (12). Following standard practices in concept drift detection, once CSS exceeds a predefined threshold ($\text{CSS}>2.3$), corresponding to the product likelihood falling below $0.1$, previously collected data are discarded. Note that we do not benchmark the cumulative Bayesian Surprise because unlike Shannon Surprise, the KL divergence nature of Bayesian Surprise makes its triggering threshold ill-defined, prohibiting the design of a reaction mechanism.
@@
Shannon and Bayesian Surprise are computed following the procedure described in Section 4.1. For MIS calculations, we discretize the range of pollution values observed in the data into $100$ bins to estimate entropy, and we set the triggering probability at $\rho=0.1$.
@@
Across all comparisons, the baseline strategies display considerable volatility. In contrast, MISRP-governed counterparts produce smoother and consistently lower error curves, highlighting the stabilizing effect of MIS through its ability to facilitate adaptive responses in dynamic environments. CSS on the other hand, is constantly triggered, leaving very few observations for estimation, thus resulting in much higher estimation error and even higher volatility than the baseline strategies.
Table 3 presents the average estimation errors and their corresponding standard errors. The standard error is measured across $10$ Monte Carlo simulations and $450$ frames. Across all sampling strategies, incorporating the MIS reaction policy yields a substantial reduction in both mean estimation error and variability. Improvements in estimation error (compared to baseline strategies) range from $24\%$ to $76\%$, while reductions in standard error range from $36\%$ to $90\%$. On the other hand, CSS shows considerable performance degradation even compared to baseline strategies, highlighting the potential negative impact of traditional concept drift detection policy in this inherently dynamic process. The quantitative results in Table 3 highlight the impact of MISRP on learning performance, demonstrating substantial improvements in estimation accuracy and stability through data stream control.
@@
Table 3: Comparison of pollution map estimation errors: baseline versus MISRP versus CSS.
Sampling Strategy Estimation Error (Baseline) Estimation Error (CSS) Estimation Error (MISRP) Mean Improvement (Over Baseline) Std. Error Improvement (Over Baseline)
SR with Shannon $6.64\pm 0.436$ $19.98\pm 0.645$ $\mathbf{1.60\pm 0.043}$ $76\%$ $90\%$
SR with Bayesian $2.79\pm 0.096$ $14.78\pm 0.374$ $\mathbf{0.87\pm 0.016}$ $69\%$ $83\%$
SC/E $2.02\pm 0.071$ $75.63\pm 2.371$ $\mathbf{1.53\pm 0.045}$ $24\%$ $36\%$
GS/QBC $2.07\pm 0.071$ $35.67\pm 1.205$ $\mathbf{1.49\pm 0.039}$ $28\%$ $45\%$
@@
So far we demonstrated that governing basic sampling strategies with MISRP can substantially enhance learning performance in dynamic environments. To provide a clearer view of how MISRP operates over time, we conduct an additional simulation examining its actions throughout the process.
In this experiment, we simulate a two-phase pollution map evolution governed by the same PDE used in earlier simulation. During the first phase (time 0250250), three pollution sources emit high levels of pollutants, and the map evolves under diffusion, decay, and wind effects. At time step $250$, the emission sources are removed, and the decay factor is reduced to one-twentieth of its original value. The system then continues evolving for an additional $50$ steps.
When the pollution sources exists and are emitting (the dynamic phase, time 0250250), the underlying process is a non-stationary process in which we expect frequent MIS triggering. When the pollution sources are gone (the stationary phase, time 251251-$300$), the pollutants in the area will eventually diffuse to a stationary existence, during which time MIS is expected to stop being triggered.
Figure 9: A visualization of estimation error progression with MISRP action overlays.
Figure 9 shows the estimation error progression with action overlays under surprise-reactive sampling based on Shannon surprise. Recall from Section 3.4 that there are two actions employed in MISRP governance: sampling adjustments and process forking. These two actions are marked as red vertical lines and green shaded regions in the plot, respectively. For clarity, we present the $20$-frame moving average of estimation error, whereas the non-smooth version is provided in the Appendix. Actions are displayed $20$ steps in advance, corresponding to their first observable effect on the smoothed error trajectory.
Several key observations emerge from the figure. First, both sampling adjustments and process forking occur frequently during the dynamic phase as expected, highlighting the effectiveness of MISRP’s action design in maintaining low estimation error. Second, sudden spikes in estimation error (circled) under MISRP governance are almost always followed by corrective actions that prevent further error growth, resulting in non-smooth error progressions after intervention. By contrast, the baseline sampling strategy allows estimation error to rise unchecked. Then, once the system enters the stationary phase, MISRP ceases intervention, aligning with the intuition that a balanced sampling strategy in a well-regulated system should not trigger Mutual Information Surprise.